# A Generalized Control Function Approach to Production Function Estimation

### v2

@@ Abstract
We generalize the control function approach to production function estimation. Our generalization accommodates scenarios in which productivity evolves jointly with other unobservable factors such as latent demand shocks and the invertibility assumption underpinning the traditional proxy variable approach fails. We provide conditions under which the output elasticity of the variable input—and hence the markup—is nonparametrically point-identified. A Neyman orthogonal moment condition ensures oracle efficiency of our GMM estimator. A Monte Carlo exercise shows a large bias for the traditional proxy variable approach that decreases rapidly and nearly vanishes for our generalized control function approach.
The proxy variable approach to production function estimation following ? (henceforth, OP) rests on two key assumptions. First, the invertibility assumption requires that there exists a function $\omega_{it}=\kappa(x_{it})$ that maps observables $x_{it}$ into the productivity $\omega_{it}$ of firm $i$ in period $t$. It is well understood that invertibility can fail because of unobserved demand heterogeneity or in imperfectly competitive environments with partially or fully unobserved rivals or changes in firm conduct (see ? and the references therein). In this paper, we denote these unobservable factors collectively as $\delta_{it}$. Besides latent demand shocks, $\delta_{it}$ may capture unobserved variation in investment opportunities or financial constraints across firms or time.
Second, the proxy variable approach rests on the assumption that productivity is governed by a Markov process with law of motion $\omega_{it}=g(\omega_{it-1})+\xi_{it}$, where $g(\omega_{it-1})=\text{E}\left[\omega_{it}|\omega_{it-1}\right]$ and the productivity innovation $\xi_{it}$ is outside the firm’s information set in period t1t-1. Replacing the autonomous Markov process with a controlled Markov process is straightforward as long as the control is observed (see ? for learning-by-exporting and ? for R&D).
@@
In this paper, we allow productivity $\omega_{it}$ and any other unobservables $\delta_{it}$ to evolve jointly and generalize the law of motion for productivity from $\omega_{it}=g(\omega_{it-1})+\xi_{it}$ to $\omega_{it}=g(\omega_{it-1},\delta_{it-1})+\xi_{it}$. This accommodates scenarios in which productivity and demand are linked through product innovations, foreign market access, or the firm’s R&D and advertising decisions. For example, a firm may move downmarket with products that are cheaper to produce but appeal to a broader set of consumers or it may move upmarket with products that are more costly to produce but appeal to more quality-conscious consumers.
One approach to dealing with additional unobservables is to extend the univariate inversion in OP to a multivariate inversion. Ensuring the existence of a function $(\omega_{it},\delta_{it})=\kappa(x_{it})$ that recovers all unobservables, however, requires making specific assumptions on the firm’s profit maximization problem (see ? for labor-augmenting productivity and ? for unobserved intermediate input prices).
@@
In this paper, we pursue another approach that does not require a multivariate inversion. We focus on estimating the production function $f(k_{it},v_{it})$, where $k_{it}$ and $v_{it}$ are inputs, and make no attempt to identify or estimate the law of motion for productivity. We provide conditions under which the output elasticity of the variable input $v_{it}$—and hence the markup following ?—is nonparametrically point-identified and can be estimated by solving a straightforward GMM problem.
We start by showing that the proxy variable approach to production function estimation following OP can be seen as controlling for “enough” of the variation in productivity to facilitate finding instruments relative to a direct instrumental variables approach. We then generalize this control function approach by adding variables to the control function. We show that this allows handling empirically relevant settings that are outside the scope of the proxy variable approach. By adding variables to the control function, our approach entails a tradeoff between increasing robustness on the one hand and decreasing efficiency and identification power on the other hand.
@@ 1 Setup
Our setup follows ? (henceforth, DL), except that we generalize the law of motion for productivity to $\omega_{it}=\text{E}\left[\omega_{it}|\omega_{it-1},\delta_{it-1}\right]+\xi_{it}=g(\omega_{it-1},\delta_{it-1})+\xi_{it}$. For concreteness we think of $\delta_{it}$ as latent demand shocks. Firm $i$ in period $t$ uses inputs $k_{it}$ and $v_{it}$ to produce output $q_{it}$ according to the production function $q_{it}=f(k_{it},v_{it})+\omega_{it}+\varepsilon_{it}$, where lower case letters denote logs. Capital $k_{it}$ is a predetermined input that is chosen in period t1t-1 whereas $v_{it}$ is freely variable and decided on in period $t$ after the firm observes $\omega_{it}$ and $\delta_{it}$. The disturbance $\varepsilon_{it}$ sits between the firm’s output $q_{it}$ as recorded in the data and the output $q^{*}_{it}=q_{it}-\varepsilon_{it}=f(k_{it},v_{it})+\omega_{it}$ that the firm planned on when it decided on the variable input $v_{it}$. It can be interpreted alternatively as measurement error or as the untransmitted component of productivity.
We assume $\text{E}\left[\varepsilon_{it}|x_{it}\right]=0$ for observables $x_{it}=\left(k_{it},v_{it},\ldots\right)$ and $\text{E}\left[\xi_{it}+\varepsilon_{it}|z_{it}\right]=0$ for instruments $z_{it}=\left(k_{it},k_{it-1},v_{it-1},\ldots\right)$. The invertibility assumption in OP ensures $\text{E}[\omega_{it-1}|x_{it-1}]=\omega_{it-1}$ or, equivalently, $\text{E}[q_{it-1}|x_{it-1}]=f(k_{it-1},v_{it-1})+\omega_{it-1}=q^{*}_{it-1}$. The timing and Markov process assumptions in OP further ensure $\text{E}[\xi_{it}+\varepsilon_{it}|z_{it},\omega_{it-1}]=0$ and thus $\text{E}\left[\xi_{it}+\varepsilon_{it}|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0$. This, in turn, ensures that $\text{E}[q_{it-1}|x_{it-1}]$ can be used as an additional instrument. We refer the reader to DL, OP, ? (henceforth, LP), and ? (henceforth, ACF) for a justification of these assumptions and further details on the setup.
@@
We write $z_{it}\setminus z^{\dagger}_{it}$ for the elements of the vector $z_{it}$ that are not contained in the vector $z^{\dagger}_{it}$. To avoid clutter, all equalities involving random variables and conditional expectations are understood to hold almost surely. Proofs are deferred to the Supplemental Appendix.
@@ 2 Results
The proxy variable approach to production function estimation following OP can be seen as a control function approach. In particular, the control function $h\left(k_{it-1},v_{it-1},\text{E}\left[q_{it-1}|x_{it-1}\right]\right)=g\left(\text{E}\left[q_{it-1}|x_{it-1}\right]-f(k_{it-1},v_{it-1})\right)$ ensures that the moment condition
$\text{E}\left[\left.\omega_{it}+\varepsilon_{it}-h\left(k_{it-1},v_{it-1},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0,$ (1)
where $\omega_{it}+\varepsilon_{it}=q_{it}-f(k_{it},v_{it})$, holds at the true production function.
Moment condition (1) follows because the law of motion for productivity is $\omega_{it}=g(\omega_{it-1})+\xi_{it}$ in the OP/LP/ACF framework and the invertibility assumption ensures $\text{E}[q_{it-1}|x_{it-1}]=f(k_{it-1},v_{it-1})+\omega_{it-1}$. Hence, we have
$\text{E}\left[\omega_{it}+\varepsilon_{it}-g(\text{E}[q_{it-1}|x_{it-1}]-f(k_{it-1},v_{it-1}))|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$
$=\text{E}\left[\omega_{it}+\varepsilon_{it}-g(\omega_{it-1})|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=\text{E}\left[\xi_{it}+\varepsilon_{it}|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0.$
Following DL, moment condition (1) can alternatively be derived without invertibility if we choose $x_{it-1}=z_{it}$. Without invertibility, however, the control function $h\left(k_{it-1},v_{it-1},\text{E}\left[q_{it-1}|x_{it-1}\right]\right)$ generally no longer equals the law of motion $g\left(\text{E}\left[q_{it-1}|x_{it-1}\right]-f(k_{it-1},v_{it-1})\right)$.
As our notation for the control function $h\left(k_{it-1},v_{it-1},\text{E}\left[q_{it-1}|x_{it-1}\right]\right)$ emphasizes, moment condition (1) requires that given $\left(k_{it-1},v_{it-1},\text{E}[q_{it-1}|x_{it-1}]\right)$ all other variables in $z_{it}$ are uncorrelated with the sum of the transmitted and untransmitted components of productivity $\omega_{it}+\varepsilon_{it}$. Moment condition (1) is less demanding than requiring that $\text{E}\left[\omega_{it}+\varepsilon_{it}|z_{it}\right]=0$ for instruments $z_{it}$ in a direct instrumental variables approach to production function estimation. It is widely understood that finding instruments in the latter approach is difficult, or perhaps even impossible, in practice (?). A key insight of OP is that all that remains of current productivity $\omega_{it}$ after controlling for lagged productivity $\omega_{it-1}$ via the control function $h\left(k_{it-1},v_{it-1},\text{E}[q_{it-1}|x_{it-1}]\right)$ is the productivity innovation $\xi_{it}$. Because $\xi_{it}$ is an independent shock, this facilitates finding instruments.
Moment condition (1) rules out some empirically relevant cases, including the joint evolution of productivity $\omega_{it}$ and latent demand shocks $\delta_{it}$. [^1] DL show that invertibility can fail in the presence of latent demand shocks. This, in turn, invalidates current capital $k_{it}$ as an instrument even if productivity $\omega_{it}$ evolves as an $AR(1)$ process separately from latent demand shocks $\delta_{it}$. DL show how to proceed without invertibility under the maintained assumption that $\omega_{it}$ evolves separately from $\delta_{it}$. To accommodate these cases and $\omega_{it}=g(\omega_{it-1},\delta_{it-1})+\xi_{it}$ as the law of motion for productivity, we generalize the control function approach in OP. To this end, we choose the special instrument $z^{\dagger}_{it}\subseteq z_{it}$ and replace moment condition (1) with the following assumption:
@@
$\text{E}\left[\left.\omega_{it}+\varepsilon_{it}-h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0$ for the control function $h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\text{E}\left[\omega_{it}+\varepsilon_{it}|z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$.
The special instrument $z^{\dagger}_{it}$ can encompass one or several components of the instruments $z_{it}$. Assumption 1 requires that given $\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$, all other variables in $z_{it}$ are uncorrelated with the sum of the transmitted and untransmitted components of productivity $\omega_{it}+\varepsilon_{it}$. The restrictiveness of Assumption 1 depends on the choice of $z^{\dagger}_{it}$. If we choose $z^{\dagger}_{it}=z_{it}\setminus(k_{it-1},v_{it-1})$, then Assumption 1 coincides with moment condition (1). If we instead choose $z^{\dagger}_{it}\subsetneq z_{it}\setminus(k_{it-1},v_{it-1})$, then Assumption 1 is less demanding: as $z^{\dagger}_{it}$ contains fewer variables, the conditioning set $z_{it}\setminus z^{\dagger}_{it}$ contains more variables and thus controls for more of the variation in productivity.
@@
As discussed above, Assumption 1 is satisfied in the OP/LP/ACF framework. More importantly, it enables going beyond this framework by not relying on invertibility and by shifting the focus from the law of motion for productivity to specifying a model for the special instrument $z^{\dagger}_{it}$. As the researcher chooses the special instrument, one can leverage institutional features or auxiliary data to justify Assumption 1. Example 1 illustrates this point.
@@
Assume that the price of the variable input $p^{V}_{it}$ evolves according to $p^{V}_{it}=\kappa(p^{V}_{it-1},\eta_{it},\tau_{i})$ for some function $\kappa$, where $\eta_{it}$ is a firm- and time-specific input price shock and $\tau_{i}$ is a firm-specific shifter such as the type of inputs the firm uses. The shifter $\tau_{i}$ may be correlated with $\delta_{it-1}$ in the law of motion for productivity. This accommodates scenarios in which the type of inputs links productivity and demand.
@@
Choosing $z_{it}=\left(k_{it},k_{it-1},v_{it-1},p^{V}_{it},\left(p^{V}_{it^{\prime}}\right)_{t^{\prime}<t}\right)$, $x_{it-1}=z_{it}$, and $z^{\dagger}_{it}=p^{V}_{it}$, we have
@@
and Assumption 1 holds. Furthermore, because $\text{E}[q_{it-1}|x_{it-1}]$ depends only on $z_{it}\setminus z^{\dagger}_{it}$, the control function simplifies from $h(z_{it}\setminus z^{\dagger}_{it}$, $\text{E}[q_{it-1}|x_{it-1}])$ to $h(z_{it}\setminus z^{\dagger}_{it})$.
@@
$\text{E}\left[\left.q_{it}-f(k_{it},v_{it})-h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0$ (3)
@@
We are interested in the output elasticity $\frac{\partial f(k_{it},v_{it})}{\partial v_{it}}$ of the variable input $v_{it}$ as the key to estimating the markup. The choice of the special instrument $z^{\dagger}_{it}$ entails a tradeoff. As noted above, as $z^{\dagger}_{it}$ contains fewer variables, the conditioning set $z_{it}\setminus z^{\dagger}_{it}$ contains more variables and thus controls for more of the variation in productivity. This leaves less exogenous variation in $z^{\dagger}_{it}$ for identifying $\frac{\partial f(k_{it},v_{it})}{\partial v_{it}}$. To ensure that $z^{\dagger}_{it}$ has sufficient identification power, we make the following assumption:
@@
Let $k_{it}\in z_{it}\setminus z^{\dagger}_{it}$. Conditional on $\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$, $z^{\dagger}_{it}$ is a complete instrument for $v_{it}$.
Assumption 2 places restrictions on the underlying economic model. In particular, it rules out the case where the law of motion for productivity is $\omega_{it}=g(\omega_{it-1})+\xi_{it}$ and the variable input demand $v_{it}=\kappa(k_{it},\omega_{it})$ is some function $\kappa$ that depends only on capital and productivity and can be inverted for productivity. [^3] Assumption 2 therefore also rules out the nonidentification result in ?. In this case, for any choice of the special instrument $z^{\dagger}_{it}\subseteq z_{it}\setminus(k_{it},k_{it-1},v_{it-1})$, $z^{\dagger}_{it}$ is independent of $\omega_{it}$ conditional on $\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ because $\omega_{it-1}$ is pinned down by $\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ and the productivity innovation $\xi_{it}$ is an independent shock. Consequently, $z^{\dagger}_{it}$ cannot generate any variation in $v_{it}=\kappa(k_{it},\omega_{it})$ after controlling for $\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$. Note that this lack-of-variation argument breaks down if latent demand shocks $\delta_{it}$ enter the law of motion for productivity or the variable input demand. Hence, latent demand shocks $\delta_{it}$ make room for Assumption 2 to hold.
@@
$\text{E}\left[\varphi(z_{it})\left(q_{it}-f(k_{it},v_{it})-h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right)\right]=0$ (4)
[…]
@@
$\tilde{\varphi}_{it}=\text{E}\left[\varphi(z_{it})\Big|z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right],\quad\tilde{q}_{it}-\tilde{f}_{it}=\text{E}\left[q_{it}-f(k_{it},v_{it})\Big|z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right],$
holds at the true production function. Moreover, if a production function satisfies moment condition (5), then it satisfies moment condition (4) for the control function $h(z_{it}\setminus z^{\dagger}_{it}$, $\text{E}[q_{it-1}|x_{it-1}])=\tilde{q}_{it}-\tilde{f}_{it}$. Finally, if $\varphi(z_{it})$ and $q_{it}-f(k_{it},v_{it})$ have finite $L^{2}$ norms, then moment condition (5) is Neyman orthogonal with respect to $L^{2}$-integrable perturbations of $\left(\tilde{\varphi}_{it},\tilde{q}_{it}-\tilde{f}_{it}\right)$.

### v3

# Ulrich Doraszelski

@@ Abstract
We develop a generalized control function approach to production function estimation. Our approach accommodates settings in which productivity evolves jointly with other unobservable factors such as latent demand shocks and the invertibility assumption underpinning the traditional proxy variable approach fails. We provide conditions under which the output elasticity of the variable input—and hence the markup—is nonparametrically point-identified. A Neyman orthogonal moment condition ensures oracle efficiency of our GMM estimator. A Monte Carlo exercise shows a large bias for the traditional approach that decreases rapidly and nearly vanishes for our generalized control function approach.
@@
As the difference of revenue and cost, a firm’s profit depends on the demand that it faces and on its production technology. Demand and cost are linked through product innovations, foreign market access, and the firm’s advertising decisions. For example, a firm may move downmarket with products that are cheaper to produce but appeal to a broader set of consumers or it may move upmarket with products that are more costly to produce but appeal to more quality-conscious foreign consumers. Cost also depends directly on the firm’s R&D decisions that, in turn, may be subject to financial constraints.
In this paper, we estimate the production function that describes the firm’s production technology. To capture the interrelation of demand and cost, we allow the productivity $\omega_{it}$ of firm $i$ in period $t$ to evolve jointly with an unobservable factor $\delta_{it}$. Besides latent demand shocks, $\delta_{it}$ may capture unobserved variation in investment opportunities or financial constraints across firms or time.
Our setting poses two difficulties for the traditional proxy variable approach to production function estimation developed by ?, ?, and ? (henceforth, OP, LP, and ACF). First, the OP/LP/ACF framework is underpinned by an invertibility assumption that in particular requires that there is no other unobservable, such as our $\delta_{it}$, besides productivity $\omega_{it}$ (the scalar unobservable assumption in ACF). Second, this framework assumes that productivity is governed by a Markov process with law of motion $\omega_{it}=g(\omega_{it-1})+\xi_{it}$. In contrast, we generalize the law of motion for productivity to $\omega_{it}=g(\omega_{it-1},\delta_{it-1})+\xi_{it}$ to allow $\omega_{it}$ and $\delta_{it}$ to evolve jointly. [^1] Replacing the autonomous Markov process in the OP/LP/ACF framework with a controlled Markov process is straightforward if the control is observed (see ? for learning-by-exporting and ? for R&D) but not if the control is unobserved.
@@
To tackle these difficulties, we develop a generalized control function approach to production function estimation. We show that our approach generalizes the control function already present in the OP/LP/ACF framework. We focus on estimating the production function and make no attempt to identify or estimate the law of motion for productivity. We provide conditions under which the output elasticity of the variable input $v_{it}$—and hence the markup following ?—is nonparametrically point-identified and can be estimated by solving a straightforward GMM problem. A Neyman orthogonal moment condition ensures oracle efficiency of our GMM estimator.
@@ 1 Setup
Our setup follows ? (henceforth, DL), except that we generalize the law of motion for productivity from $\omega_{it}=g(\omega_{it-1})+\xi_{it}$ to $\omega_{it}=g(\omega_{it-1},\delta_{it-1})+\xi_{it}$, where $g(\omega_{it-1},\delta_{it-1})=\text{E}\left[\omega_{it}|\omega_{it-1},\delta_{it-1}\right]$ and $\xi_{it}$ is the productivity innovation. For concreteness we think of $\delta_{it}$ as latent demand shocks. Firm $i$ in period $t$ uses inputs $k_{it}$ and $v_{it}$ to produce output $q_{it}$ according to the production function
$q_{it}=f(k_{it},v_{it})+\omega_{it}+\varepsilon_{it},$
where lower case letters denote logs. Capital $k_{it}$ is a predetermined input that is chosen in period t1t-1 whereas $v_{it}$ is freely variable and decided on in period $t$ after the firm observes $\omega_{it}$ and $\delta_{it}$. The disturbance $\varepsilon_{it}$ sits between the firm’s output $q_{it}$ as recorded in the data and the output $q^{*}_{it}=q_{it}-\varepsilon_{it}=f(k_{it},v_{it})+\omega_{it}$ that the firm planned on when it decided on the variable input $v_{it}$. It can be interpreted alternatively as measurement error or as the untransmitted component of productivity.
In what follows, we let $x_{it}=\left(k_{it},v_{it},\ldots\right)$ denote a vector of observables and $z_{it}=\left(k_{it},k_{it-1},v_{it-1},\ldots\right)$ a vector of instruments. As in DL, these vectors can contain additional variables besides the ones listed.
In the OP/LP/ACF framework, estimation proceeds in two steps. The first step assumes $\text{E}[\varepsilon_{it}|x_{it}]=0$ and flexibly or nonparametrically estimates the conditional expectation $\text{E}[q_{it}|x_{it}]$. The second step assumes $\text{E}[\xi_{it}+\varepsilon_{it}|z_{it}]=0$ and estimates the production function and the law of motion for productivity using GMM.
@@
For any two vectors $a$ and $b$, we write $a\setminus b$ for the elements of $a$ that are not contained in $b$. To avoid clutter, equalities involving random variables and conditional expectations are understood to hold almost surely. Proofs are deferred to the Supplemental Appendix.
@@ 2 Results
We choose the special instrument $z^{s}_{it}\subseteq z_{it}$ and partition $z_{it}$ as $z_{it}=(z^{s}_{it},z^{c}_{it})$. In the following assumption, $z^{c}_{it}$ serve as control variables whereas the special instrument $z^{s}_{it}$ is excluded from the control function.
@@
$\text{E}\left[\left.\omega_{it}+\varepsilon_{it}-h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0$ for the control function $h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$.
The special instrument $z^{s}_{it}$ can consist of one or several components of $z_{it}$. As our notation for the control function $h\left(z^{c}_{it},\text{E}\left[q_{it-1}|x_{it-1}\right]\right)$ emphasizes, Assumption 1 requires that conditional on $\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$, the special instrument $z^{s}_{it}$ is uncorrelated with the sum of the transmitted and untransmitted components of productivity $\omega_{it}+\varepsilon_{it}$. In contrast, Assumption 1 allows the control variables $z^{c}_{it}$ to be correlated with productivity.
Because it controls for a part of the variation in productivity, Assumption 1 is less demanding than the condition $\text{E}[\omega_{it}+\varepsilon_{it}|z_{it},\text{E}[q_{it-1}|x_{it-1}]]=0$ used in a conventional instrumental variables approach to production function estimation. It is widely understood that finding instruments in the latter approach is difficult, or perhaps even impossible, in practice (?).
We further illustrate Assumption 1 with two examples. Example 1 shows that Assumption 1 holds in the OP/LP/ACF framework.
Example 1 (OP/LP/ACF framework).
In the OP/LP/ACF framework, the law of motion for productivity is $\omega_{it}=g(\omega_{it-1})+\xi_{it}$. The GMM estimation in the second step uses the moment condition [^2] In the OP/LP/ACF framework, $\text{E}[q_{it-1}|x_{it-1}]$ can be used as an additional instrument. Because the invertibility assumption ensures $\omega_{it-1}=\text{E}\left[q_{it-1}|x_{it-1}\right]-f(k_{it-1},v_{it-1})$ and the timing and Markov process assumptions further ensure $\text{E}[\xi_{it}+\varepsilon_{it}|z_{it},\omega_{it-1}]=0$, we have $\text{E}\left[\xi_{it}+\varepsilon_{it}|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0$.
$\text{E}\left[\left.q_{it}-f(k_{it},v_{it})-g\left(\text{E}\left[q_{it-1}|x_{it-1}\right]-f(k_{it-1},v_{it-1})\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0.$ (1)
@@
Because $q_{it}-f(k_{it},v_{it})=\omega_{it}+\varepsilon_{it}$, moment condition (1) implies Assumption 1 if we choose $z^{c}_{it}=\left(k_{it-1},v_{it-1}\right)$, $z^{s}_{it}=z_{it}\setminus z^{c}_{it}$, and $h\left(z^{c}_{it},\text{E}\left[q_{it-1}|x_{it-1}\right]\right)=g\left(\text{E}\left[q_{it-1}|x_{it-1}\right]-f(k_{it-1},v_{it-1})\right)$.
Note that in the OP/LP/ACF framework the invertibility assumption ensures that lagged productivity $\omega_{it-1}=\text{E}\left[q_{it-1}|x_{it-1}\right]-f(k_{it-1},v_{it-1})$ can be recovered from observables. Moment condition (1) and Assumption 1 are therefore closely related. The difference is that moment condition (1) restricts the control function to take on the specific form implied by the Markov process and invertibility assumptions while Assumption 1 leaves it unrestricted.
Example 1 clarifies that the OP/LP/ACF framework can be seen as a control function approach. A key insight of OP is that all that remains of current productivity $\omega_{it}$ after controlling for lagged productivity $\omega_{it-1}$ via the control function $h\left(z^{c}_{it},\text{E}\left[q_{it-1}|x_{it-1}\right]\right)=g\left(\text{E}\left[q_{it-1}|x_{it-1}\right]-f(k_{it-1},v_{it-1})\right)$ is the productivity innovation $\xi_{it}$. Because $\xi_{it}$ is an independent shock, this facilitates finding instruments.
Following DL, Assumption 1 can also be implied without the invertibility assumption if we choose $x_{it-1}=z_{it}$ in addition to $z^{c}_{it}=\left(k_{it-1},v_{it-1}\right)$ and $z^{s}_{it}=z_{it}\setminus z^{c}_{it}$. Without invertibility, however, the control function $h\left(z^{c}_{it},\text{E}\left[q_{it-1}|x_{it-1}\right]\right)$ generally no longer equals the law of motion $g\left(\text{E}\left[q_{it-1}|x_{it-1}\right]-f(k_{it-1},v_{it-1})\right)$. DL maintain that the law of motion for productivity is $\omega_{it}=g(\omega_{it-1})+\xi_{it}$.
Relative to the OP/LP/ACF framework and DL, Assumption 1 generalizes the control function approach to production function estimation by allowing us to add variables to the control function. As the conditioning set $\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ contains more variables and thus controls for a larger part of the variation in productivity, the special instrument $z^{s}_{it}$ contains fewer variables. Adding variables to the control function therefore makes Assumption 1 less demanding. This increases robustness.
@@
Perhaps even more importantly, Assumption 1 allows us to go beyond the OP/LP/ACF framework by not relying on invertibility and accommodating the joint evolution of productivity $\omega_{it}$ and latent demand shocks $\delta_{it}$. It shifts the focus from the law of motion for productivity to specifying a model for the special instrument $z^{s}_{it}$. As the researcher chooses the special instrument, one can leverage institutional features or auxiliary data to justify Assumption 1. Example 2 illustrates this point.
@@
Assume that the price of the variable input $p^{V}_{it}$ evolves according to $p^{V}_{it}=\kappa(p^{V}_{it-1},\eta_{it},\tau_{i})$ for some function $\kappa$, where $\eta_{it}$ is a firm- and time-specific input price shock and $\tau_{i}$ is a firm-specific shifter such as the type of inputs the firm uses. The shifter $\tau_{i}$ may be correlated with $\delta_{it-1}$ in the law of motion for productivity. This accommodates settings in which the type of inputs links productivity and demand.
@@
Choosing $z_{it}=\left(k_{it},k_{it-1},v_{it-1},p^{V}_{it},\left(p^{V}_{it^{\prime}}\right)_{t^{\prime}<t}\right)$, $x_{it-1}=z_{it}$, $z^{s}_{it}=p^{V}_{it}$, and $z^{c}_{it}=z_{it}\setminus z^{s}_{it}$, we have
@@
and Assumption 1 holds for the control function $h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\text{E}\left[\omega_{it}+\varepsilon_{it}|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right]$. Furthermore, because $\text{E}[q_{it-1}|x_{it-1}]$ depends only on $z^{c}_{it}$, the control function simplifies from $h(z^{c}_{it}$, $\text{E}[q_{it-1}|x_{it-1}])$ to $h(z^{c}_{it})$.
@@
$\text{E}\left[\left.q_{it}-f(k_{it},v_{it})-h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right|z_{it},\text{E}[q_{it-1}|x_{it-1}]\right]=0$ (3)
@@
We are interested in the output elasticity $\frac{\partial f(k_{it},v_{it})}{\partial v_{it}}$ of the variable input $v_{it}$ as the key to estimating the markup. The choice of the special instrument $z^{s}_{it}$ entails a tradeoff. As noted above, as the conditioning set $\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ contains more variables and thus controls for a larger part of the variation in productivity, the special instrument $z^{s}_{it}$ contains fewer variables. This leaves less exogenous variation in $z^{s}_{it}$ for identifying $\frac{\partial f(k_{it},v_{it})}{\partial v_{it}}$. To ensure that $z^{s}_{it}$ has sufficient identification power, we make the following assumption:
@@
Let $k_{it}\in z^{c}_{it}$. Conditional on $\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$, $z^{s}_{it}$ is a complete instrument for $v_{it}$.
Assumption 2 places restrictions on the underlying economic model. In particular, it rules out the case where the law of motion for productivity is $\omega_{it}=g(\omega_{it-1})+\xi_{it}$ and the variable input demand $v_{it}=\kappa(k_{it},\omega_{it})$ is some function $\kappa$ that depends only on capital and productivity and can be inverted for productivity. [^4] Assumption 2 therefore also rules out the nonidentification result in ?. In this case, for any choice of the special instrument $z^{s}_{it}\subseteq z_{it}\setminus(k_{it},k_{it-1},v_{it-1})$, $z^{s}_{it}$ is independent of $\omega_{it}$ conditional on $\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ because $\omega_{it-1}$ is pinned down by $\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ and the productivity innovation $\xi_{it}$ is an independent shock. Consequently, $z^{s}_{it}$ cannot generate any variation in $v_{it}=\kappa(k_{it},\omega_{it})$ after controlling for $\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$. Note that this lack-of-variation argument breaks down if latent demand shocks $\delta_{it}$ enter the law of motion for productivity or the variable input demand. Hence, latent demand shocks $\delta_{it}$ make room for Assumption 2 to hold.
@@
$\text{E}\left[\varphi(z_{it})\left(q_{it}-f(k_{it},v_{it})-h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)\right)\right]=0$ (4)
holds at the true production function. GMM estimation based on moment condition (4) can be conducted by viewing the control function $h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)$ as a nuisance parameter and estimating it alongside the production function, as in the OP/LP/ACF framework. [^5] The law of motion for productivity cannot generally be recovered from estimates of the production and control functions. However, depending on the number of control variables $z^{c}_{it}$, the dimension of the control function may be high. This increases the asymptotic variance of the estimates and can create numerical challenges for minimizing the GMM objective function.
@@
$\tilde{\varphi}_{it}=\text{E}\left[\varphi(z_{it})|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right],\quad\tilde{q}_{it}-\tilde{f}_{it}=\text{E}\left[q_{it}-f(k_{it},v_{it})|z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right],$
holds at the true production function. Moreover, if a production function satisfies moment condition (5), then it satisfies moment condition (4) for the control function $h\left(z^{c}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\tilde{q}_{it}-\tilde{f}_{it}$. Finally, if $\varphi(z_{it})$ and $q_{it}-f(k_{it},v_{it})$ have finite $L^{2}$ norms, then moment condition (5) is Neyman orthogonal with respect to $L^{2}$-integrable perturbations of $\left(\tilde{\varphi}_{it},\tilde{q}_{it}-\tilde{f}_{it}\right)$.